A girl went to the market with 10 dollars. Spent 4 dollars on meat, half of what was left on fish and one-third of what was then left on vegetables. How much had he left?
Answer by Cromlix(4381)

Hi there,
10 dollars
10 - 4 for meat
= 6 dollars
Half of this on fish
1/2 of 6 = 3
6 -3 = 3
3 dollars left
one third of 3 on vegetables
1/3 x 3 = 1
3 - 1 = 2
2 dollars left.
Hope this helps :-)